OpenStudy (anonymous):

How to integrate v^2 (1-v) ^6 let u=1-v then i get : (1-u)^2(u^6) what to do next?

OpenStudy (jhannybean):

\[\large \int\limits v^2(1-v)^6dv\]is that your question?

OpenStudy (jhannybean):

\[\large {u=1-v \ \\v= 1-u}\]

OpenStudy (jhannybean):

\[\large dv = -du\]

OpenStudy (jhannybean):

\[\large - \int\limits [(1-u)^2\cdot u^6]du\]

OpenStudy (anonymous):

Expand (1-u)^2 then multiply u^6 inside the expanded terms . Then do as usual integration .

OpenStudy (anonymous):

if in case of the (1-u) has very big index?

OpenStudy (anonymous):

Dont forget -ve sign

OpenStudy (jhannybean):

What does expanding (1-u)^2 give you, first?

OpenStudy (anonymous):

1-2u +u^2 then all of the term times by u^6

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

but if i have v^8 * (v-1)^6 ?

OpenStudy (jhannybean):

\[\large - \int\limits\limits [(1-2u+u^2)\cdot u^6]du\]

OpenStudy (anonymous):

Like ??? u have made substitution to reduce the index of (1-v)^6 to something u had already known

OpenStudy (anonymous):

i mean 1-v --- in the baracket

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Then expand [(V-1)^3]^2

OpenStudy (jhannybean):

Now,what do you get when you multiply u^6 into all these terms?

OpenStudy (anonymous):

i get it, thanks guys

OpenStudy (jhannybean):

Ok :)

OpenStudy (anonymous):

It becomes lengthy but ultimately u"ll be able to solve it . Just use what u already knew :)

OpenStudy (anonymous):

you could also integrate by parts using a tabular method

OpenStudy (anonymous):

for efficiency: http://people.whitman.edu/~hundledr/courses/M244S07/IntByParts.pdf